OpenStudy (rissyroo118):

Solve, using the substitution method. y – 3x = 7 21 + 9x = 3y (Points : 1) There is no solution. There are an infinite number of solutions. The solution is (1, 10) The solution is (28, 0) help

jagr2713 (jagr2713):

send screen shot of question

OpenStudy (rissyroo118):

y – 3x = 7 21 + 9x = 3y

jagr2713 (jagr2713):

x = 3 , y = 2

jagr2713 (jagr2713):

wait no

OpenStudy (rissyroo118):

lol , take your time. i have another one like this problem there really confusing to me

OpenStudy (kirbykirby):

y – 3x = 7 y = 7 + 3x Plug in this y in the 2nd equation: 21 + 9x = 3(7 + 3x) 21 + 9x = 21 + 9x 21 - 21 = 9x - 9x 0 = 0 When you encounter this situation, the 2nd equation is just a multiple of the 1st equation, and is therefore the same line. So there are infinitely many solutions (both lines are lying on top of each other)